OpenStudy (anonymous):

what is the equation in standard form of the line which passes through (-2,6) and has a slope of -1?

OpenStudy (anonymous):

use this formula: y-y1=m(x-x1) (x1,y1)=(-2,6) m=-1

OpenStudy (anonymous):

y-y1 = m(x-x1) y-6=-(x+2) y-6=-x-2 y = -x+4

OpenStudy (anonymous):

nice job